4. Solve y - y - 12y = 144.+ 12.5 with the conditions y(0) = 5 and y(0) = -0.5..
Answer #1

Sol % - Given, Second order ordinery differential equation, 7--12y = 144.23 +12.5 - and conditions, Yo = 5 YOO = -0.5 Now,Now, To find Yps. of 12.5 i. Jas = (D2-D-12) 14412 . (D2-D-107 (D2-9-12) X3 (12.5) eo* we know, OPP. 292 so fofcan provided f- Jos = -12[23+4+* +34 +32 + tex] - 12:6 = -12 x3 – X-37-38-120,5 12 Now. A So The Gmplete solution. - y = dox + y p. I 0 = c